(no subject)

Carlos Escobar (escobar@fsgi02)
Wed, 1 Sep 1999 07:04:12 -0500 (CDT)

Shift Summary for Owl shift, 01/Sept/99

On Shift: Val, Carlos

Hours of beam available at all intensities:7 hrs
Hours of beam available at nominal intensity:7 hrs
Hours of E832 data (special runs not included):3.5 hrs
Hours of special runs (laser scans, muon runs...):1 hr

Accounting of lost time:
Discovering that NM2S3 trip was the reason for th scalers alarms: 1.5 hours.
Muon run initialization problems and KTRIG problems: 1 hr.

========================================================================
Runs ended this shift:
Run 14386 - 13.4 M events - BEAM_832

Tapes brought to FCC this shift: QK*369-371, QKB129.
(including tapes for online split)

========================================================================
Accesses during shift: NONE

========================================================================
Overview of shift:
Smooth E832 running until the mid of the shift, A tripped magnet (NM2S3)
spoiled the rest of the shift.
MUON run from that point onwards. Problems with KQ.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14386 00:00 E832 data taking. Run started on previous shift.

01:05 ktev4a: 91% full. Deleted /ktev4a/data/VETO/14293*data
/14309*data
/14344*data


127 02:38 High intensity spill (>7E12) causes livetime=0.
Restarted triggers. All is well.


02:41 During monitoring of the histograms it was noticed
that:
- SA2 has two channels with relatively low ADC values.
Channels 7 and 16. SET notified.
- ET3 threshold around 28.09 GeV.
- Cosmic T17 pedestal disagrees with database by
30 counts.

159 03:20 KTEVANA error, stream 5, unknown crate, KFREND

03:23 KTEV2A full. Deleted /data/VETO/*14309*dat
*14344*dat

03:55 For the last 20 or so spills (168-190) we have had
scalers warnings repeatedly:SA3/REG_TOP =0.37 [0.2,0.3]
MU2_1HT/REG_TOP =0.70 [0.2,0.4]

04:08 Discover that NM2S3 has tripped off. Last 30 spills may
not be usable (spill 168 onwards).
RUN 14386 stopped. Tapes full.

04:15 RUN 14387 STARTED
Tapes QK*372-374,QKB129.
Waiting for MCR to bring NM2S3 up.

04:50 Two MCR operators stop by to inform us that they
can not ramp NM2S3 up.

Decide to take a MUON RUN.

STOP RUN 14387. No useful data.

05:00 Run 14388 started.
Init failed. Hangs. Do a drt_fresh start.
Magnets set ok. MCR calls to say NM2S2 has been set
appropriatelly. Beam stop in.

05:30 Problems with KQ initialization.

06:00 Reset KQ's. drt_fresh start again.
Initialized run 14391
Start run 14391.
Trigger errors. Killed KTRIG process and restarted it
as KFREND does not help. This fixes the problems.
Muon Run has several KTEVANA errors and we are dead
every third spill or so.

07:00 Hand over to day crew.

========================================================================
Follow up Notes on Previous Shift

o get flaky power supply from hall and return to prep (Ed took care of it)

========================================================================
Items to Follow up on Next Shift

o do a laser scan at the start of shutdown.

o page Rick Kessler after laser scan.

o fill in oil for the vacuum pump (camera 13).